Twice a number is decreased by 9, then multiplied by 4. The result is 8 less than 10 times the number. What is the number?
Answer by John10(297)

4(2x - 9)= 10x - 8
8x - 36 = 10x - 8
10x - 8x = -36 + 8
2x = -28
x = -14
John10:)